OpenStudy (anonymous):

Which of the following best describes the amount of gas used over time while driving 55 miles per hour on a highway?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Not sure what the choices given are represented as but you can check the answers by creating the ratio of miles to gas used: \[ 55 \frac{ miles }{ hour }\] "Gas used over time": If we say its gas per hour then we have: \[\frac{ gas }{ hour }\] But if we further say we use gallons of gas then we have: \[\frac{ gallons }{ hour }\] It all depends on the choices given

OpenStudy (anonymous):

discrete and non-linear continuous and non-linear discrete and linear continuous and linear